# Neural network-driven domain decomposition for efficient solutions to the Helmholtz equation

### v1

@@ 1 Introduction
where $\mathcal{D}$ is some differential operator, $u$ is the unknown solution and $\mathbf{x}$ is the space variable, with $\Omega$ being the space domain. Note that (1) does not contain a boundary operator; this is intentional as boundary condition will be handled by PMLs. To solve (1), we use Physics-Informed Neural Networks (PINNs) proposed in pinns . PINNs train a Neural Network (NN) to solve a differential equation by directly approximating the solution, i.e., defining $v(\cdot,\boldsymbol{\theta})\approx u(\cdot)$, where $\boldsymbol{\theta}$ represents a set of trainable parameters of the neural network. The main advantages of PINNs are their mesh-free nature and the fact that we end up with the approximation of a function rather than a discretised solution, unlike conventional numerical methods. Based on pinns , the following loss function is minimized to train the PINN
$L(\theta)$ $=\frac{1}{N}\sum_{i=1}^{N_{I}}\underbrace{\left(\mathcal{D}[v(x_{i},\theta)]-f(x_{i})\right)^{2}}_{\text{PDE residual}},$ (2)
where $\{x_{i}\}_{i=1}^{N_{I}}$ is a set of collocation points sampled in the domain. As mentioned above, we do not include a boundary loss term here, as the PML formulation ensures wave absorption at the boundaries, removing the necessity for explicit boundary constraints as in standard PINNs.
We move from the standard PINNs formulation to its domain-decomposed extension, namely the Finite Basis PINNs (FBPINNs) introduced in  fbpinns , which decompose the domain into overlapping subdomains. We want to combine them with PML to simulate free-space wave propagation. The goal of this work is to assess the performance of the combination of FBPINNs and PML approach for solving the Helmholtz equation.
@@ 2.1 Perfectly Matched Layer
To approximate the Sommerfeld radiation condition, the problem is reformulated using a PML transformation introduced by Bérenger pml_basic in 1994. The PML offers a more robust approximation than e.g. Robin boundary conditions by creating a finite, artificial absorbing layer around the simulation area.
@@
We are looking for the solution in the region $\Omega=[-L,L]^{2}$, considering a symmetrical domain without loss of generality. To apply the PML, the computational domain $\Omega=[-L-L_{\text{PML}},L+L_{\text{PML}}]^{2}$, where $L_{\text{PML}}$ represents the width of our absorbing layer. The computational domain is illustrated in Figure 1 (left), where the blue square corresponds to the physical region and the red layer represents the PML. At this stage, we consider a single computational domain. The implementation of a PML uses a coordinate stretching technique, thus, we define the coordinate stretching functions
@@
$\sigma(x)=\sigma_{0}k\begin{cases}\Bigl(\dfrac{-L-x}{L_{\text{PML}}}\Bigr)^{2},&x<-L,\\[6.0pt] 0,&-L\leq x\leq L,\\[6.0pt] \Bigl(\dfrac{x-L}{L_{\text{PML}}}\Bigr)^{2},&x>L,\end{cases}$
The tuning constant $\sigma_{0}$ and the PML thickness are specified parameters in the numerical experiments. The function $\sigma$ is defined following pml_dd ; pml_ref , but other choices are also possible. Then, thanks to the absorbing layer, we no longer impose a boundary condition explicitly, since it is incorporated into the equation itself. As a result, the differential operator with PML $\mathcal{D}_{\text{PML}}$, associated with the Helmholtz equation, defined as
@@ 2.2 FBPINNs
This section defines the mathematical formulation of FBPINNs according to fbpinns . First, we split the global domain $\Omega$ into $J$ overlapping subdomains $\smash{\{\Omega_{j}\}_{j\in\{1,\dots,J\}}}$. To achieve smoother and more accurate solutions, these subdomains are constructed with an overlap $\eta$, which will be specified in the numerical experiments. This overlap is introduced to avoid dealing explicitly with continuity constraints at subdomain interfaces. The approximate solution to our problem (1) is then defined as
$\tilde{u}(\mathbf{x},\boldsymbol{\theta})=\sum_{j=1}^{J}\varphi_{j}(\mathbf{x})\,v_{j}(\mathbf{x};\theta_{j}),$ (5)
where $\tilde{u}$ is the weighted combination of the local neural networks $v_{j}$, each trained on its corresponding subdomain $\Omega_{j}$. The window function $\varphi_{j}$ restricts the influence of $v_{j}$ to its local subdomain while ensuring a smooth transition between neighbouring regions.
Each window function must be smooth, differentiable, and form a partition of unity, i.e. $\smash{\sum_{j=1}^{J}\varphi_{j}(\mathbf{x})=1}$. Like in multilevel , we use a cosine-based window function
@@
Here, $\mu_{j,n}$ denotes the centre of the $j$-th subdomain along the $n$-th spatial axis, and $\smash{\psi_{j,n}=\nicefrac{{\eta\Delta_{n}}}{{2}}}$ defines its half-width, with $\Delta_{n}$ the subdomain size along dimension $n$.
@@
$L(\boldsymbol{\theta})=\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{D}_{\text{PML}}[\tilde{u}(x_{i},\boldsymbol{\theta})]-g(x_{i})\right)^{2}.$ (6)
@@ 3 Numerical Results
Some PINNs and FBPINNs settings and hyperparameters are fixed across all experiments. Specifically, we use a learning rate of 0.001, a standart multilayer perceptron, 5000 collocation points and sine activation function. We divide the domain into four subdomains when using FBPINNs.
We compare the Adam + L-BFGS optimization technique with the Energy Natural Gradient Descent (ENGD) method. Adam and L-BFGS are state-of-the-art optimizers for training PINNs, commonly used in the literature due to their reliability and convergence properties. ENGD introduced in NG modifies the standard gradient update as follows
$\theta_{t+1}=\theta_{t}-\,G_{E}^{-1}(\theta_{t})\,\nabla_{\theta}L(\theta_{t}),$
where $G_{E}(\theta)=\nabla_{\theta}L(\theta)^{\top}\nabla_{\theta}L(\theta)$ is the energy matrix derived from the physical energy functional
$L(\theta)=E(u_{\theta})=\int_{\Omega}\left(\mathcal{D}u-g\right)^{2}\,\mathrm{d}x.$
This approach takes into account the shape of the loss surface when updating the parameters. For experiments with ENGD and Adam optimizer in combination with L-BFGS, where L-BFGS is initiated at a training ratio of 0.7, we use the architecture 3 hidden layers with 32 neurons each for PINNs and 3 hidden layers with 16 neurons in every layer for each NN of FBPINNs.
@@
From the Table 1 we can see that both methods perform similarly at low frequencies ($k=0.57$) for both methods of optimizations. However, as the wavenumber increases, the standard PINNs quickly deteriorate, showing errors close to unity already for $k=4.51$, while the FBPINNs maintain a slightly better accuracy. We can observe a similar situation with the imaginary part from Table 2, but by comparing the results, we can conclude that despite our assistance to the imaginary part in the form of preliminary training with another source term, it is still approximated worse than the real part. We assume that this effect appears because there is a local minimum where the imaginary part tends to vanish, which makes the optimization harder for this part. The difference between PINNs and FBPINNs is clearly visible in the convergence curves in Figure 2, where the FBPINNs show a faster decrease in loss and achieve a lower final value, especially for high frequencies. Nevertheless, when the wight of PML is too thin (for example $L_{\text{PML}}=0.25\lambda$), both approaches struggle to converge due to insufficient wave absorption at the boundaries.
@@
We also compare the performance of the standard Adam + L-BFGS optimizer with the ENGD.
At low frequencies, both optimizers provide similar results, although ENGD demonstrating faster stabilization of the loss that we can see from Figure 2. As the frequency increases, the difference becomes clearer: while the standard optimizer struggles to converge, ENGD succeeds in maintaining stable training and correctly reproducing the oscillatory solution. From the real and imaginary part approximations shown in Figure 3, we observed that ENGD captures the fine wave structures missed by Adam + L-BFGS, even in the imaginary part. Regarding the errors reported in Tables 1 and 2, the results are most evident when we move to the frequency of $k=4.5$, since in this case, especially for the imaginary part, Adam + L-BFGS gives the worst error, while ENGD achieve an accurate approximation. The training speed of ENGD is about six times lower than that of Adam + L-BFGS for both PINNs and FBPINNs. This is mainly due to the additional computation of the energy matrix at each iteration, which significantly increases the computational cost. The use of double precision arithmetic further contributes to the slowdown. In practice, switching to hardware that handles double precision more efficiently could reduce this gap, although the method would still remain slower overall.
@@ 4 Conclusion
In this work, we showed that FBPINNs combined with PML boundaries can effectively handle wave absorption and provide more accurate solutions than standard PINNs. While the improvement remains moderate, the obtained results confirm that the approach is effective and has potential for further refinement. The use of natural gradient preconditioning also helped to stabilize training and achieve better convergence. In future work, we plan to refine the combination of FBPINNs and PML by introducing internal absorbing layers and to extend the framework toward multilevel FBPINNs as in multilevel for more complex wave problems (higher frequencies, heterogeneous media, $\dots$).

### v2

@@ 1 Introduction
where $\mathcal{D}$ is some differential operator, $u$ is the unknown solution and $\mathbf{x}$ is the space variable, with $\Omega$ being the space domain. Note that (1) does not contain a boundary operator; this is intentional as boundary condition will be handled by PMLs. To solve (1), we use Physics-Informed Neural Networks (PINNs) proposed in Raissi et al. (2019). PINNs train a Neural Network (NN) to solve a differential equation by directly approximating the solution, i.e., defining $v(\cdot,\boldsymbol{\theta})\approx u(\cdot)$, where $\boldsymbol{\theta}$ represents a set of $P$ trainable parameters of the neural network. The main advantages of PINNs are their mesh-free nature and the fact that we end up with a continuous functional approximation of the solution, defined over the entire domain, rather than a discretised solution on a fixed mesh as in conventional numerical methods. Based on Raissi et al. (2019), the following loss function is minimized to train the PINN
$L(\boldsymbol{\theta})$ $=\frac{1}{N}\sum_{i=1}^{N_{I}}\underbrace{\big(\mathcal{D}[v_{\boldsymbol{\theta}}](\mathbf{x}_{i})-g(\mathbf{x}_{i})\big)^{2}}_{\text{PDE residual}},$ (2)
where $\{\mathbf{x}_{i}\}_{i=1}^{N_{I}}$ is a set of collocation points sampled in the domain, and where we have defined $v_{\boldsymbol{\theta}}=\mathbf{x}\mapsto v(\mathbf{x},\boldsymbol{\theta})$. As mentioned above, we do not include a boundary loss term here, as the PML formulation ensures wave absorption at the boundaries, removing the necessity for explicit boundary constraints as in standard PINNs.
We move from the standard PINNs formulation to its domain-decomposed extension, namely the Finite Basis PINNs (FBPINNs) introduced in  Moseley et al. (2023), which decomposes the domain into overlapping subdomains. In this paper, we combine them with PML to simulate free-space wave propagation. The goal of this work is thus to assess the performance of the combination of FBPINNs and PMLs for solving the Helmholtz equation.
@@ 2.1 Perfectly Matched Layer
To approximate the Sommerfeld radiation condition, the problem is reformulated using a PML transformation introduced by Bérenger Berenger (1994) in 1994. The PML offers a more robust approximation than e.g. Robin boundary conditions by creating a finite, artificial absorbing layer around the simulation area.
@@
We are looking for the solution in the region $\Omega=[-L,L]^{2}$, considering a symmetrical rectangular domain without loss of generality, since equivalent results can be obtained for non-symmetric rectangular domains. While the method can be adapted to more general domain geometries by modifying the functions defined below, such extensions are not considered in this work for the sake of conciseness. To apply the PML, the computational domain $\Omega=[-L-L_{\text{PML}},L+L_{\text{PML}}]^{2}$, where $L_{\text{PML}}$ represents the width of our absorbing layer. What do you think about the correction The computational domain is illustrated in Figure 1 (left), where the blue square corresponds to the physical region and the red layer represents the PML. At this stage, we consider a single computational domain. The implementation of a PML uses a coordinate stretching technique, thus, we define the coordinate stretching functions
@@
$\sigma(x)=\sigma_{0}k\left(\dfrac{\max\!\left(|x|-L,\,0\right)}{L_{\text{PML}}}\right)^{2}$
The tuning constant $\sigma_{0}$ and the PML thickness are specified parameters in the numerical experiments. The function $\sigma$ is defined following Bootland et al. (2024); Song et al. (2021), but other choices are also possible. Then, thanks to the absorbing layer, we no longer impose a boundary condition explicitly, since it is incorporated into the equation itself. As a result, the differential operator with PML $\mathcal{D}_{\text{PML}}$, associated with the Helmholtz equation, defined as
@@ 2.2 FBPINNs
This section defines the mathematical formulation of FBPINNs according to Moseley et al. (2023). First, we split the global domain $\Omega$ into $J$ overlapping subdomains $\smash{\{\Omega_{j}\}_{j\in\{1,\dots,J\}}}$. To achieve smoother and more accurate solutions, these subdomains are constructed with an overlap $\eta$, which will be specified in the numerical experiments. This overlap is introduced to avoid dealing explicitly with continuity constraints at subdomain interfaces. The approximate solution to our problem (1) is then defined as
$\tilde{u}(\mathbf{x},\boldsymbol{\theta})=\sum_{j=1}^{J}\varphi_{j}(\mathbf{x})\,v_{j}(\mathbf{x};\boldsymbol{\theta}_{j}),$ (5)
where $\tilde{u}$ is the weighted combination of the local neural networks $v_{j}$, each trained on its corresponding subdomain $\Omega_{j}$. The window function $\varphi_{j}$ restricts the influence of $v_{j}$ to its local subdomain while ensuring a smooth transition between neighboring regions.
Each window function must be smooth, differentiable, and form a partition of unity, i.e. $\smash{\sum_{j=1}^{J}\varphi_{j}(\mathbf{x})=1}$. Like in Dolean et al. (2024), we use a cosine-based window function
@@
Here, $\mu_{j,n}$ denotes the center of the $j$-th subdomain along the $n$-th spatial axis, and $\smash{\psi_{j,n}=\nicefrac{{\eta\Delta_{n}}}{{2}}}$ defines its half-width, with $\Delta_{n}$ the subdomain size along dimension $n$.
@@
$L(\boldsymbol{\theta})=\frac{1}{N}\sum_{i=1}^{N}\big(\mathcal{D}_{\text{PML}}[\tilde{u}_{\boldsymbol{\theta}}](\mathbf{x}_{i})-g(\mathbf{x}_{i})\big)^{2},$ (6)
where $\tilde{u}_{\boldsymbol{\theta}}=\mathbf{x}\mapsto\tilde{u}(\mathbf{x},\boldsymbol{\theta})$.
@@ 3 Numerical Results
Some PINNs and FBPINNs settings and hyperparameters are fixed across all experiments. Specifically, we use a learning rate of 0.001, a standard multilayer perceptron, 5000 collocation points and sine activation function. We divide the domain into four subdomains when using FBPINNs.
We compare the Adam + L-BFGS optimization technique with the Energy Natural Gradient Descent (ENGD) method. Adam and L-BFGS are state-of-the-art optimizers for training PINNs, commonly used in the literature due to their reliability and convergence properties. ENGD, introduced in Müller and Zeinhofer (2023), arises by projecting the optimal function-space descent direction onto the tangent space of the neural-network manifold, yielding a preconditioned gradient in parameter space, that best approximates this optimal direction. It modifies the standard gradient update as follows:
$\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\,G^{\dagger}(\boldsymbol{\theta}_{t})\,\nabla_{\theta}L(\boldsymbol{\theta}_{t}),$
where $G^{\dagger}$ is the Moore-Penrose pseudoinverse of the matrix $G$ with components
$\big(G(\boldsymbol{\theta})\big)_{ij}=\int_{\Omega}\,\partial_{\theta_{i}}\!\big(\mathcal{D}_{\text{PML}}[\tilde{u}_{\boldsymbol{\theta}}]\big)\!(\mathbf{x})\ \partial_{\theta_{j}}\!\big(\mathcal{D}_{\text{PML}}[\tilde{u}_{\boldsymbol{\theta}}]\big)\!(\mathbf{x})\ d\mathbf{x}$
for all $(i,j)\in\{1,\dots,P\}$. For experiments with ENGD and Adam optimizer in combination with L-BFGS, where L-BFGS is initiated at a training ratio of 0.7, we use the architecture 3 hidden layers with 32 neurons each for PINNs and 3 hidden layers with 16 neurons in every layer for each NN of FBPINNs.
@@
From Table 1 we can see that both methods perform similarly at low frequencies ($k=0.57$) for both methods of optimizations. However, as the wavenumber increases, the standard PINNs quickly deteriorate, showing errors close to unity already for $k=4.51$, while the FBPINNs maintain a slightly better accuracy. We can observe a similar situation with the imaginary part from Table 2, but by comparing the results, we can conclude that despite our assistance to the imaginary part in the form of preliminary training with another source term, it is still approximated worse than the real part. We assume that this effect appears because there is a local minimum where the imaginary part tends to vanish, which makes the optimization harder for this part. The difference between PINNs and FBPINNs is clearly visible in the convergence curves in Figure 2, where the FBPINNs show a faster decrease in loss and achieve a lower final value, especially for high frequencies. Nevertheless, when the wight of PML is too thin (for example $L_{\text{PML}}=0.25\lambda$), both approaches struggle to converge due to insufficient wave absorption at the boundaries.
@@
We also compare the performance of Adam + L-BFGS to ENGD. At low frequencies, both optimizers provide similar results, although ENGD demonstrating faster stabilization of the loss that we can see from Figure 2. As the frequency increases, the difference becomes clearer: while the standard optimizer struggles to converge, ENGD succeeds in maintaining stable training and correctly reproducing the oscillatory solution. From the real and imaginary part approximations shown in Figure 3, we observed that ENGD captures the fine wave structures missed by Adam + L-BFGS, even in the imaginary part. Regarding the errors reported in Tables 1 and 2, the results are most evident when we move to the frequency of $k=4.5$, since in this case, especially for the imaginary part, Adam + L-BFGS gives the worst error, while ENGD achieve an accurate approximation. The training speed of ENGD is about six times lower than that of Adam + L-BFGS for both PINNs and FBPINNs. This is mainly due to the additional computation of the energy matrix at each iteration, which significantly increases the computational cost. The use of double precision arithmetic further contributes to the slowdown. In practice, switching to hardware that handles double precision more efficiently could reduce this gap, although the method would still remain slower overall.
@@ 4 Conclusion
In this work, we showed that FBPINNs combined with PML boundaries can effectively handle wave absorption and provide more accurate solutions than standard PINNs. While the improvement remains moderate, the obtained results confirm that the approach is effective and has potential for further refinement. The use of natural gradient preconditioning also helped to stabilize training and achieve better convergence. In future work, we plan to refine the combination of FBPINNs and PML by introducing internal absorbing layers and to extend the framework toward multilevel FBPINNs as in Dolean et al. (2024) for more complex wave problems (higher frequencies, heterogeneous media, $\dots$).